OpenStudy (zzr0ck3r):

Show that \(l\) is a continuouse linear functional on \(C([0,1]) \\ \). \(l(f)=\int_0^1f(t)h(t)dt\) where \(h\in\mathscr{L^1}([0,1])\)

OpenStudy (zzr0ck3r):

I already showed that \(|||l|||\le ||f||_{[0,1]}||h||_1\) where \(||h||_1\)is the max norm; I'm not sure if that's standard notation. I assume that this is the norm of the linear functional, but none of the methods we were playing with worked for me so far. I took it to my teacher and he said that I should not worry about it, and that he should not have assigned it. But I have played with it to long to let it go.

OpenStudy (zzr0ck3r):

@eliassaab

OpenStudy (zzr0ck3r):

if its a pain in the but don't worry about it.

OpenStudy (anonymous):

\[ |l(f)|=\left|\int_0^1f(t)h(t)dt\right|\le\int_0^1|f(t)h(t)|dt\le \sup_{0\le t \le 1 }|f(t)|\int_0^1| h(t)|dt=||f||_{C[0,1]} ||h||_1 \] That is what you did. That is enough to show that l is contnious. If \( f_n \to f \) in C[0,1] then\[ |l(f_n ) -l(f)|=|l(f_n-f)| \le ||f_n-f||_{C[o,1]} ||h||_1 \] You are done

OpenStudy (anonymous):

A linear operator T is continuous if and only if \[ ||T(x)|| \le M ||x||\]

OpenStudy (anonymous):

Where M is positive constant

OpenStudy (zzr0ck3r):

I'm confused on how this shows what the norm is.